OpenStudy (lilday12):

Which choices are in the solution set of the equation below? Check all that apply. 15x = 60 A. 4 B. 3 C. 3.5 D. E. 0.25 F. 4.5

OpenStudy (unklerhaukus):

divide both side of the equation by 15, what do you get?

OpenStudy (lilday12):

a and e

OpenStudy (unklerhaukus):

?, how do you get two options?

OpenStudy (lilday12):

i did 15/60=0.25 and 60/15=4

OpenStudy (unklerhaukus):

15x = 60 15x / 15 = 60 /15 x = 4

OpenStudy (lilday12):

it A

OpenStudy (unklerhaukus):

yesh